OpenStudy (anonymous):

if the difference of a number and four is tripled, and the result is eight more than the number. find the number.

OpenStudy (anonymous):

STFU

OpenStudy (kc_kennylau):

Come on, we keep doing work for you but you can't learn anything, what's the point?

OpenStudy (anonymous):

you helped me with one problem and barely explained it.

OpenStudy (kc_kennylau):

How would you express "the difference a number and four"?

OpenStudy (kc_kennylau):

I'm trying to break down the question into pieces so that you can follow.

OpenStudy (anonymous):

idk x-4?

OpenStudy (kc_kennylau):

yes, you're getting it :)

OpenStudy (kc_kennylau):

How would you express "eight more than the number"?

OpenStudy (anonymous):

would the equation be +8?

OpenStudy (kc_kennylau):

Try it again :) ok, you've declared the number to be x, than how would you express a number which is "eight more than the number", or "eight more than x"?

OpenStudy (anonymous):

sorry I didn't mean to say the equation /.\

OpenStudy (kc_kennylau):

But that's still not correct, try it again :)

OpenStudy (anonymous):

or try and use the numbers and use your logic on what numbers you could plug in and subtract four and then the anwser - x equals 8

OpenStudy (anonymous):

that is correct logic right kc_kennylau

OpenStudy (kc_kennylau):

yep

OpenStudy (phi):

start with difference of a number and four is tripled, "the difference a number and four" changes to x - 4 now we have x-4 is tripled to figure this part out, think what does 2 is tripled mean ? it means 2*3=6 or 6 is tripled? it means 6*3=18. In other words, tripled means "multiply by 3" x-4 is tripled becomes (x-4)*3 (notice we put x-4 in parens because we multiply the whole thing by 3.

OpenStudy (phi):

you started with difference of a number and four is tripled, and the result is eight more than the number and so far have 3(x-4) is eight more than the number is becomes equal sign. eight becomes 8 so that means 3(x-4) is eight more than the number becomes 3(x-4) = 8 more than the number can you finish changing the words into math?

OpenStudy (anonymous):

if you are talking to me phi as changing my words to math i will next time thanks for the feedback